5. (1 mark) Laurens Volleyball Company produces two types of volleyba Both products pass through two producing departments.
Answer #1

5)

A. Calculation of activity rate for machine-related costs based on machine hours

Total machine-related costs = $240,000

Total machine hours = Machine hours for Official game + Machine hours for Volley lite

                                  = 10,000 + 40,000

                                  = 50,000

Activity rate = Total machine-related costs / Total machine hours

                   = $240,000 / 50,000

                   = $4.8 per machine hour

B. Calculation of activity rate for other manufacturing overhead costs based on packing orders

Total other manufacturing overhead costs = $240,000

Total packing orders = Packing orders for Official game + Packing orders for Volley lite

                                 = 30,000 + 30,000

                                 = 60,000

Activity rate = Total other manufacturing overhead costs / Total packing orders

                    = $240,000 / 60,000

                    = $4 per packing order

Both machine-related costs and other manufacturing overhead costs form part of Manufacturing Overhead costs.

C. Calculation of Manufacturing Overhead cost per unit for Official game

With the given data, Manufacturing Overhead cost per unit for Official game

Manufacturing Overhead cost per unit
Particulars Amount
A Machine related costs $48,000
(10,000 machine hours * $4.8 per machine hour)
B Other manufacturing overheads $120,000
(30,000 packing orders * $4 per packing order)
C Total Manufacturing Overheads $168,000
D No. of units produced 20,000
E Manufacturing Overhead cost per unit {E = (C/D)} $8.4

D. Calculation of Manufacturing Overhead cost per unit for Volley lite

With the given data, Manufacturing Overhead cost per unit for Volley lite

Manufacturing Overhead cost per unit
Particulars Amount
A Machine related costs $192,000
(40,000 machine hours * $4.8 per machine hour)
B Other manufacturing overheads $120,000
(30,000 packing orders * $4 per packing order)
C Total Manufacturing Overheads $312,000
D No. of units produced 200000
E Manufacturing Overhead cost per unit {E = (C/D)} $1.56